OpenStudy (anonymous):

A system of equations is shown below: 6x - 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if I sum both equations, I get: 9x=9 so what is x?

OpenStudy (michele_laino):

@FEARLESS_JOCEY

OpenStudy (anonymous):

1 ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

nw we substitute x=1, for example, into the first equation, so we get: 6-5y=5 5y=1 so what is y?

OpenStudy (michele_laino):

@FEARLESS_JOCEY

OpenStudy (anonymous):

y = 5 ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

no, please y=1/5

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

So what is the next step?

OpenStudy (michele_laino):

sorry again: we have 5y=1 so y=1/5 then the solution of your system is:x=1, and y=1/5

OpenStudy (michele_laino):

then the requested x-coordinate is 1

OpenStudy (anonymous):

Wow! You helped a lot, thanks!

OpenStudy (michele_laino):

thanks! :)